OpenStudy (anonymous):

Solve for x 64(x^2) = 25

OpenStudy (lalaly):

\[(8x-5)(8x+5)=0\]

OpenStudy (anonymous):

divide by 64, square root (25/64)

OpenStudy (lalaly):

\[x=\frac{5}{8}\]\[x=\frac{-5}{8}\]

OpenStudy (anonymous):

x^2 =25/64 Square root 25/64 =0.625

OpenStudy (anonymous):

I'm thinking 5/8

OpenStudy (lalaly):

u fractorise it difference between two squares (8x-5)(8x+5)=0

OpenStudy (anonymous):

if you see my above reply, it is \[\sqrt{25/64}\] which leads to \[\pm5/8\]

OpenStudy (anonymous):

+5/8, -5/8

OpenStudy (anonymous):

64 x^2-25 = 0 so, x = -5/8 or x = 5/8

OpenStudy (anonymous):

thanks..!! i am the best..##